# Curve fitting on a quantum annealer for an advanced navigation method

### v1

@@ V Summary and conclusions
Finding optimal speed profiles is an optimization problem that grows exponentially in time. Dynamic programming tames the combinatorial explosion of trajectories by solving the problem backward in time. The state-dependent optimal value function as a solution of the Bellman equation solves the problem linearly in the number of time steps but it requires a grid-based exploration of the entire state space. If the dimension of the latter is high, the well-known phenomena of “curse of dimensionality” poses a severe obstacle, particularly for large real-world use cases. An approximation of the optimal value function and thus its continuous representation mitigates the rounding error that occurs in the discrete grid-based setup.
Such approximations can often be reduced to curve fitting in one dimension. In this work, we formulated least-squares curve fitting, where we allowed for a rather general approximator in form of a finite linear combination of standardized (basis) functions, as a QUBO that is suited to be solved with quantum annealing. For simple functions/data points, we found that quantum annealing in its current state is in principle able to deliver comparable results to a classical computer. However, regardless of the choice of basis functions, curve fitting tasks that result in a QUBO size larger than about $60\times 60$ (i.e., with about 60 logical variables or more) cannot be solved accurate enough on a D-Wave annealer yet. The reason for that is the difficult low-energy landscape of the resulting Ising Hamiltonian.
@@
Nevertheless, the quantum annealing-oriented curve fitting principally works. Toward a real-word use case, it is for example possible to solve problems—albeit still in an exploratory fashion—within the topic of just-in-time navigation using that approach for approximating the value function in the Bellman equation of dynamic programming.
As a final remark, there are many other ways to approximate functions using a quantum computer. A promising approach is described in Ref. [55], which employs parameterized quantum circuits on a gate-based quantum computer using partial Fourier series. This seems to work with fewer qubits but the training process is significantly different. There are also ways to fit curves on a photonic quantum computer, see, e.g., Ref. [56]. There, the approximator is a variational quantum circuit built as an continuous-variable architecture that encodes quantum information in continuous degrees of freedom.
Acknowledgements.
We acknowledge funding and support from the German Federal Ministry for Economic Affairs and Climate Action (Bundesministerium für Wirtschaft und Klimaschutz) through the PlanQK initiative. This project (HA project no. 1362/22-67) is financed with funds of LOEWE—Landes-Offensive zur Entwicklung Wissenschaftlich-ökonomischer Exzellenz, Förderlinie 3: KMU-Verbundvorhaben (State Offensive for the Development of Scientific and Economic Excellence).

### v2

@@ V Summary and conclusions
Finding optimal speed profiles is an optimization problem that grows exponentially in time. Dynamic programming tames the combinatorial explosion of trajectories by solving the problem backward in time. The state-dependent optimal value function as a solution of the Bellman equation solves the problem linearly in the number of time steps but it requires a grid-based exploration of the entire state space. If the dimension of the latter is high, the well-known phenomena of the curse of dimensionality poses a severe obstacle, particularly for large real-world use cases. An approximation of the optimal value function and thus its continuous representation mitigates the rounding error that occurs in the discrete grid-based setup.
Such approximations can often be reduced to curve fitting in one dimension. In this work, we formulated least-squares curve fitting, where we allowed for a rather general approximator in form of a finite linear combination of standardized (basis) functions, as a QUBO that is suited to be solved with quantum annealing. For simple functions/data points, we found that quantum annealing in its current state is in principle able to deliver comparable results to a classical computer. However, curve fitting tasks that result in a QUBO size larger than roughly about $60\times 60$ (i.e., with about 60 logical variables or more) cannot be solved accurate enough on a D-Wave annealer yet. The reason for that is the difficult low-energy landscape of the resulting Ising Hamiltonian. Moreover, QUBO problems described by global basis functions (e.g., Chebyshev polynomials) are even more difficult to solve, hence even smaller problems can have a poor solution quality.
@@
Nevertheless, the quantum annealing-oriented curve fitting principally works. Regarding a real-world use case—albeit still in an exploratory fashion—we investigated the linear-quadratic optimal control problem of just-in-time arrival of a vessel, where we approximated the value function in the Bellman equation. With an increasing number of basis functions, the solution space can be better represented, leading to more accurate value estimates compared to the grid-based method. Nevertheless, there was hope that the quantum annealing process would yield lower errors in larger search spaces. Unfortunately, we find that D-Wave’s current quantum annealing hardware does not yet appear to be capable of meeting this expectation.
Acknowledgments
We acknowledge funding and support from the German Federal Ministry for Economic Affairs and Climate Action (Bundesministerium für Wirtschaft und Klimaschutz) through the PlanQK initiative (01MK20005).
This project (HA project no. 1362/22-67) is financed with funds of LOEWE—Landes-Offensive zur Entwicklung Wissenschaftlich-ökonomischer Exzellenz, Förderlinie 3: KMU-Verbundvorhaben (State Offensive for the Development of Scientific and Economic Excellence).
Declarations
Author contributions
Conceptualization, funding acquisition, and supervision: B. Harrach and W. Mergenthaler. Formal analysis, investigation, and software: P. Isserstedt and D. Jaroszewski. Validation: P. Isserstedt, D. Jaroszewski, and F. Paul. Writing (original draft): P. Isserstedt and D. Jaroszewski. Writing (review and editing): P. Isserstedt, D. Jaroszewski, and F. Paul. This list uses the Contributor Role Taxonomy (CRediT) [71].
Competing interests
The authors have no competing interests to declare that are relevant to the content of this article.